OpenStudy (nathan917):

From a committee of 6 girls and 4 boys, a name is selected. Then another name is selected. What is the probability that both people drawn will be girls? (1 point) nine over twenty-five two-fifteenths three-fifths one-third

OpenStudy (anonymous):

Similar to last problem: (6/10) times (5/9) = 1/3 This is a "non-replacement" problem. The total probability is: P(1st is a girl) times P(2nd is a girl) And since one girl is already picked and therefore unavailable, ther are only 5 girls and 9 total people left for the second pick.

OpenStudy (nathan917):

I Got it. probability of drawing a girl in the first pick x probability of drawing a girl in the 2nd pick = 6/10 x 5/9 = (6x5) / (10x9) = 30 / 90 = 1 / 3

OpenStudy (anonymous):

You are understanding this well now!

OpenStudy (nathan917):

Yes.

OpenStudy (anonymous):

You basically did what I did and modeled after the previous problem. Good job, indeed!

OpenStudy (anonymous):

6c1/10c1 * 5c1/9c1 = 1/3